OpenStudy (anonymous):

find q if p varies directly as the square of q and k=8 and p=64

OpenStudy (abmon98):

\[p=k*\sqrt{q}\]

OpenStudy (anonymous):

thats it?

OpenStudy (abmon98):

\[64=8*\sqrt{q}\] \[64/8=\sqrt{q}\] \[(8)^2=q\]

OpenStudy (anonymous):

@Abmon98 what happened to the square root?

OpenStudy (abmon98):

When we move the square to the other side of the equation the fraction 1/2 is inverted to 2

OpenStudy (anonymous):

@Abmon98 im sorry i dont get it :(

OpenStudy (abmon98):

so you should learn that y^1/2 means the square root of y we get rid of that square root , by squaring both sides \[(\sqrt{q})^2=8^2\] \[(q ^{1/2})^2=q\] 2 gets cancelled by the other 2

OpenStudy (anonymous):

so the answer is 64 ? @Abmon98